# Euclid Quick Data Release (Q1)

### v1

@@ 6 Success rates
[…]
Using the DESI redshifts in the EDF-N as a reference, we can estimate the success rate of SPE redshift measurements. Since the DESI redshift distribution is biased towards lower redshifts and brighter objects than for the objects targeted by Euclid, the results must be taken with caution, as discussed in Sect. 3.1. A better internal assessment will be possible at DR1, when a subset of the EDS (Euclid Collaboration: Scaramella et al. 2022) will be available. The full EDS will benefit from 15 multiple Wide-like red grism (1206–1892 Å) visits, supplemented by a further 25 visits with the blue grism (900–1300 Å). This has been designed as to retrieve virtually pure and complete reference samples, at the EWS depth (thanks to larger sensitivity), minimal spectral confusion between adjacent objects (thanks to multiple observing angles), and reduced redshift degeneracies (thanks to the extended blue-plus-red wavelength range). We will therefore have a reliable and complete control sample to calibrate EWS observations – including measurements of $\mathrm{H}\,\alpha$ fluxes, which will allow a proper estimate of the completeness. In addition, realistic simulations will also be released along with DR1.
To obtain a quantitative estimate of the success rate, relative to the DESI sample, we compute the fraction of objects with $|z_{\rm SPE}-z_{\rm DESI}|/(1+z_{\rm DESI})<0.003$ in various sub-samples. This range corresponds to the nominal $3\,\sigma$ interval around SPE measurements. We focus on the range $1.60.9<z_{\rm SPE}<$ 1.6, i.e., the usual cosmological target range of Euclid, where $\mathrm{H}\,\alpha$ falls within the red grism, but limited to $z=1.6$, corresponding to the upper cut-off of the DESI sample. In the following sections, we study the dependence of the success rate as a function of the SPE redshift-probability threshold (gal_spe_z_prob). As expected, a high threshold produces a more reliable sample. However, this reduces the size of the recovered sample and thus the completeness, which cannot be estimated yet in the absence of full-depth EDFs, and the cutoff in magnitude.
@@ 6.1 Success rate for non-cosmological science
[…]
We also studied the SPE success rate outside the redshift range where $\mathrm{H}\,\alpha$ can be detected. In Table 2, we summarise the performance obtained in various redshift slices. If we consider all three classes of objects and use a 0.99 redshift probability threshold for the best class, we reach a success rate of 53% at low $z$ ($z<0.1$) without adding any additional criteria. This result is encouraging, but it falls to only 7% if we select objects classified as galaxies, and we see that the apparently good results are mainly driven by stars (see Sect. 4). Not surprisingly, the results are also poor at $0.90.1<z<$ 0.9, with an 8% success rate, explained mainly by the absence of any strong spectral features in the red-grism wavelength range. The performance is even poorer above $z=1.8$, with a 2% success rate. This result is not so surprising, since no bright spectral feature is expected to be detected in the red grism above $z=1.8$ given the sensitivity of NISP.
Our results illustrate the importance of applying appropriate selection criteria when using spectroscopic data from the Q1 release. Measurements for galaxies in the Euclid target redshift range, where $\mathrm{H}\,\alpha$ is covered by the red grism provide encouraging results. The same is true for stars. Outside of this redshift range, SPE results alone cannot be used to build a reasonably pure redshift sample for statistical studies; they must be complemented with additional data such as photometry (Cagliari et al. 2024) or morphological criteria that will be provided by the high quality Euclid images. A visual inspection of the spectra can also help to refine small samples for specific studies (Euclid collaboration: Quai et al. in prep.).
Table 2: Summary of the success rate of SPE for various measured redshift ranges (lines) and SPE redshift probability thresholds $P_{\rm SPE}$ (columns). These are derived from a comparison with DESI data, and require some care in interpretation. DESI cannot detect emission-line galaxies beyond $z=1.6$, and therefore all our correct redshifts between 1.6 and 1.8 will not be included here.
@@ 6.2 Success rate for cosmology goals
Figure 8: Success rate as a function of the redshift probability for the cosmological sample. The blue dashed line is computed using all types of objects as input and SPE is used to identify the galaxies. The orange solid line is computed using objects classified by galaxies by both SPE and DESI. The grey area represents the Euclid performance goal for cosmology.
@@
In Fig. 3, we compare the redshift obtained by SPE and by DESI for the selection described in the previous paragraph and a SPE redshift probability larger than 0.99. Since the Euclid selection for cosmology is based on the measured $\mathrm{H}\,\alpha$ line, the measured redshifts are all between 0.9 and 1.8. The results are dramatically improved compared to simpler selections discussed in Sect. 3 with most of the galaxies on the 1:1 relation. However, we also find a small population of galaxies with a lower DESI redshift, while only one object has a higher DESI redshift. Since they are not aligned, we can reasonably assume that these are spurious SPE redshifts that do not arise from simple line misidentifications. These $z<0.9$ sources erroneously measured by SPE at $z>0.9$ presumably arise from noise in the Euclid red-grism spectral range, and it is not surprising to have a small fraction of such spurious redshifts out of the large population of galaxies with $H_{\scriptscriptstyle\rm E}<22.5$ and $z<0.9$ (see Sect. 2.1).
In Fig. 8, we present the SPE success rate as a function of the redshift properties for galaxies with measured properties matching the criteria of the cosmological sample. As explained in Sect. 6.1, we consider only the $1.60.9<z<$ 1.6 range, where both DESI and SPE are reliable. If we select galaxies based on the SPE classification only, we reach an 80 % success rate for z_spe_proba0.996absent0.996\,$\geq$ 0.996≥ 0.996 (blue dashed line). However, the SPE-only classification is not highly reliable (Sect. 4), and we will eventually benefit from Euclid morphological and photometric information to discard stars and lower redshift galaxies from the sample. To obtain a performance forecast closer to the expected methods that will be used in Euclid, we computed the success rate for a clean sample classified by both SPE and DESI as galaxies. A z_spe_proba threshold of 0.985 is enough to obtain an 80% success rate, and we can reach 89% for a stringent z_spe_proba0.999absent0.999$\geq$ 0.999≥ 0.999 criterion.
In short, we have demonstrated that success rates higher than the specifications (see Table 2 for a summary) can be obtained for the cosmological sample if we select the sample carefully. For the future Euclid DR1, further adjustments will be possible in order to establish the optimal threshold that yields the best compromise between success rate and sample size, in order to maximise the performance for cosmology. These future analyses will use the EDS data, which are not part of the Q1.
@@ 7 Conclusions
We have presented the first data release (Q1) of the SPE PF results. If we properly combine the various quality indicators provided by the pipeline, we showed that we can reach a success rate above 80% in the redshift range targeted for cosmology ($1.80.9<z<$ 1.8). The redshift precision (approximately $10^{-3}$) and accuracy (better than $3\times 10^{-5}$) are also excellent. This is encouraging for the future DR1 and the first cosmological results, but the completeness will still need to be assessed when EDS data will be available and all EWS sources down to $H_{\scriptscriptstyle\rm E}=24$ instead of 22.5 will have been processed.
As would be expected, the results are less good outside of this redshift range, since no strong spectral features are then observable within the NISP red-grism spectral range, and these data should be used with caution for non-cosmology science. Future deep-field observations including the blue grism should dramatically improve Euclid’s performance in this regime. In addition, the sample of star candidates selected by SPE PF based on spectroscopy alone is 75% pure. This is very promising for the study of cool stars (Banados et al., in prep.).
Finally, the current results are only a first step and we can expect major improvements in the future. The PF-SIR products injected into SPE PF still contain important artefacts (Euclid Collaboration: Copin et al. 2025), which will be corrected in the next versions of the pipeline. This should dramatically reduce the fraction of spurious redshifts and enable us to use less stringent selection criteria, increasing the size of the high-reliability samples. The SPE PF pipeline itself will also improve. Deep-learning approaches have been developed to estimate the redshift reliability and classify the spectra. We already trained and tested such algorithms on simulations and they outperformed the current algorithm. However, we will need future real EDS data to build a proper training set for the analysis of real EWS data. Finally, the selection of galaxies for cosmological analyses will also benefit from non-spectroscopic information (e.g., morphology or photometry), allowing us to obtain sufficiently pure and complete samples.
@@ Acknowledgements.
This work has made use of the Euclid Quick Release Q1 data from the Euclid mission of the European Space Agency (ESA), 2025, https://doi.org/10.57780/esa-2853f3b. The Euclid Consortium acknowledges the European Space Agency and a number of agencies and institutes that have supported the development of Euclid, in particular the Agenzia Spaziale Italiana, the Austrian Forschungsförderungsgesellschaft funded through BMK, the Belgian Science Policy, the Canadian Euclid Consortium, the Deutsches Zentrum für Luft- und Raumfahrt, the DTU Space and the Niels Bohr Institute in Denmark, the French Centre National d’Etudes Spatiales, the Fundação para a Ciência e a Tecnologia, the Hungarian Academy of Sciences, the Ministerio de Ciencia, Innovación y Universidades, the National Aeronautics and Space Administration, the National Astronomical Observatory of Japan, the Netherlandse Onderzoekschool Voor Astronomie, the Norwegian Space Agency, the Research Council of Finland, the Romanian Space Agency, the State Secretariat for Education, Research, and Innovation (SERI) at the Swiss Space Office (SSO), and the United Kingdom Space Agency. A complete and detailed list is available on the Euclid web site (www.euclid-ec.org). This research used data obtained with the Dark Energy Spectroscopic Instrument (DESI). DESI construction and operations is managed by the Lawrence Berkeley National Laboratory. This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of High-Energy Physics, under Contract No. DEAC0205CH11231, and by the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility under the same contract. Additional support for DESI was provided by the U.S. National Science Foundation (NSF), Division of Astronomical Sciences under Contract No. AST-0950945 to the NSFs National Optical-Infrared Astronomy Research Laboratory; the Science and Technology Facilities Council of the United Kingdom; the Gordon and Betty Moore Foundation; the Heising-Simons Foundation; the French Alternative Energies and Atomic Energy Commission (CEA); the National Council of Science and Technology of Mexico (CONACYT); the Ministry of Science and Innovation of Spain (MICINN), and by the DESI Member Institutions: www.desi.lbl.gov/collaborating-institutions. The DESI collaboration is honoured to be permitted to conduct scientific research on Iolkam Duag (Kitt Peak), a mountain with particular significance to the Tohono Oodham Nation. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the U.S. National Science Foundation, the U.S. Department of Energy, or any of the listed funding agencies.

### v2

@@ 6 Success rates
Figure 7: Success rate (fraction of correct redshifts) as a function of the SPE redshift probability threshold for objects classified as galaxies and with a measured redshift between 0.9 and 1.6. The various solid coloured lines correspond to all objects (dotted blue line), objects with an $\mathrm{H}\,\alpha$ S/N above 3.5 (dash-dotted orange line), and objects with an $\mathrm{H}\,\alpha$ S/N above 5 (dashed red line). Finally, the solid purple line presents the result of the following selection: $\mathrm{H}\,\alpha$ S/N>\,>\,5, measured $\mathrm{H}\,\alpha$ flux larger than $2\times 10^{-16}\,\mathrm{erg\,s}^{-1}\,\mathrm{cm}^{-2}$, and a line width strictly smaller than the upper bound of the prior ($680\,$\mathrm{km}\text{\,}{\mathrm{s}}^{-1}$\$).
Using the DESI redshifts in the EDF-N as a reference, we can estimate the success rate of SPE redshift measurements. In this paper, we call the ‘success rate’ the fraction of objects with a correct redshift for a given selection based on the measured parameters and the quality metrics returned by SPE. [^1] We could have called it ‘purity’, but the notion of purity have a very strict definition in the Euclid context and we thus choose another terminology to avoid any ambiguity. Since the DESI redshift distribution is biased towards lower redshifts and brighter objects than the objects targeted by Euclid, the results must be taken with caution, as discussed in Sect. 3.1. We discuss the impact of the choice of DESI as reference sample in Appendix A. A better internal assessment will be possible at DR1, when a subset of the EDS (Euclid Collaboration: Scaramella et al., 2022) will be available. The full EDS will benefit from 15 multiple EWS-like red grism (1206–1892 Å) visits, supplemented by a further 25 visits with the blue grism (900–1300 Å). This has been designed to retrieve virtually pure and complete reference samples at the EWS depth (thanks to larger sensitivity), minimal spectral confusion between adjacent objects (thanks to multiple observing angles), and reduced redshift degeneracies (thanks to the extended blue-plus-red wavelength range). We will therefore have a reliable and complete control sample with which to calibrate EWS observations – including measurements of $\mathrm{H}\,\alpha$ fluxes, which will allow for a proper estimate of the completeness. In addition, realistic simulations will also be released along with DR1.
To obtain a quantitative estimate of the success rate, relative to the DESI sample, we computed the fraction of objects with $|z_{\rm SPE}-z_{\rm DESI}|/(1+z_{\rm DESI})<0.003$ in various sub-samples. This range corresponds to the nominal $3\,\sigma$ interval around SPE measurements. We focused on the range $0.9<z_{\rm SPE}<1.6$, which is the usual cosmological target range of Euclid, where $\mathrm{H}\,\alpha$ falls within the red grism, but limited to $z=1.6$, corresponding to the upper cut-off of the DESI sample. In the following sections, we study the dependence of the success rate as a function of the SPE redshift-probability threshold (gal_spe_z_prob). As was expected, a high threshold produces a more reliable sample. However, this reduces the size of the recovered sample and thus the completeness, which cannot be estimated yet in the absence of full-depth EDFs, and the cut-off in magnitude.
@@ 6.1 Success rate for non-cosmological science
The success rate for various quality selections is shown in Fig. 7. In the absence of any additional selection (blue), the reliability goes from 14% for a null SPE redshift probability threshold to 27% for a 0.99 threshold. The results can be improved dramatically by adding $\mathrm{H}\,\alpha$ S/N criteria (3.5 in orange and 5 in red). With these additional criteria, we reach 61% and 75% success rates for a 0.99 probability threshold and S/N thresholds of 3.5 and 5, respectively. Finally, we add additional criteria on the flux and the width of the $\mathrm{H}\,\alpha$ line found by template fitting. We keep only the sources with $\mathrm{H}\,\alpha$ flux larger than $2\times 10^{-16}\,\mathrm{erg\,s}^{-1}\,\mathrm{cm}^{-2}$ and also discard all objects for which the estimated line width saturates at the maximum value of the allowed range ($680\,$\mathrm{km}\text{\,}{\mathrm{s}}^{-1}$\$), which is a signal of a spurious broad feature mistaken as emission line. With this selection, we reach a success rate of 85% for a $\texttt{gal\_spe\_z\_prob}=0.99$ threshold.
We also studied the SPE success rate outside the redshift range in which $\mathrm{H}\,\alpha$ can be detected. In Table 2, we summarise the performance obtained in various redshift slices. If we consider all three classes of objects and use a 0.99 redshift probability threshold for the best class, we reach a success rate of 53% at low $z$ ($z<0.1$) without adding any additional criteria. This result is encouraging, but it falls to only 7% if we select objects classified as galaxies and we see that the apparently good results are mainly driven by stars (see Sect. 4). Not surprisingly, the results are also poor at $0.1<z<0.9$, with an 8% success rate, explained mainly by the absence of any strong spectral features in the red-grism wavelength range. The performance is even poorer above $z=1.8$, with a 2% success rate. This result is not so surprising, since no bright spectral feature is expected to be detected in the red grism above $z=1.8$ given the sensitivity of NISP.
Our results illustrate the importance of applying appropriate selection criteria when using spectroscopic data from the Q1 release. Measurements for galaxies in the Euclid target redshift range, where $\mathrm{H}\,\alpha$ is covered by the red grism provide encouraging results. The same is true for stars. Outside of this redshift range, SPE results alone cannot be used to build a reasonably pure redshift sample for statistical studies; they must be complemented with additional data such as photometry (Cagliari et al., 2024) or morphological criteria that will be provided by the high-quality Euclid images. A visual inspection of the spectra can also help to refine small samples for specific studies (Euclid collaboration: Quai et al. in prep.).
Table 2: Summary of SPE success rate  [^2] The success rate is given for various measured redshift ranges (lines) and SPE redshift probability thresholds, $P_{\rm SPE}$ (columns). These rates are derived from a comparison with DESI data, and some care is required when interpreting them. DESI cannot detect emission-line galaxies beyond $z=1.6$, and therefore all our correct redshifts between 1.6 and 1.8 will not be included here.
@@ 6.2 Success rate for cosmology goals
Figure 8: Success rate (fraction of correct redshifts) as a function of the redshift probability for the cosmological sample. The dashed blue line is computed using all types of objects as input and SPE is used to identify the galaxies. The solid orange line is computed using objects classified by galaxies by both SPE and DESI. The grey area represents the Euclid performance goal for cosmology.
The measurement of reliable redshifts between $z=0.9$ and $z=1.8$ is the first crucial step towards Euclid’s cosmological goals based on galaxy clustering measurements. Although we do not have a true reference sample to test the cosmological performance with, we can nevertheless draw useful indications by properly trimming the Q1 and DESI samples. We thus selected a galaxy sample similar to the baseline cosmological sample: $\mathrm{H}\,\alpha$ flux larger than $2\times 10^{-16}\,\mathrm{erg\,s}^{-1}\,\mathrm{cm}^{-2}$ and $\textrm{S/N}>3.5$. In addition, we kept only objects for which the line width is narrower than the prior limit of $680\,$\mathrm{km}\text{\,}{\mathrm{s}}^{-1}$\$ to exclude artefacts (Sect. 3). The size limit of 0.5 is not included in this study and even better results could be obtained by adding this criterion.
In Fig. 3, we compare the redshift obtained by SPE and by DESI for the selection described in the previous paragraph and a SPE redshift probability larger than 0.99. Since the Euclid selection for cosmology is based on the measured $\mathrm{H}\,\alpha$ line, the measured redshifts are all between 0.9 and 1.8. The results are dramatically improved compared to simpler selections discussed in Sect. 3 with most of the galaxies on the 1:1 relation. However, we also find a small population of galaxies with a lower DESI redshift, while only one object has a higher DESI redshift. Since they are not aligned, we can reasonably assume that these are spurious SPE redshifts that do not arise from simple line misidentifications. These $z<0.9$ sources erroneously measured by SPE at $z>0.9$ presumably arise from noise in the Euclid red-grism spectral range and it is not surprising to have a small fraction of such spurious redshifts out of the large population of galaxies with $H_{\scriptscriptstyle\rm E}<22.5$ and $z<0.9$ (see Sect. 2.1).
@@
In short, we have demonstrated that success rates higher than the specifications (see Table 2 for a summary) can be obtained for the cosmological sample if we select the sample carefully. For the future Euclid DR1, further adjustments will be possible in order to establish the optimal threshold that yields the best compromise between success rate and sample size, in order to maximise the performance for cosmology. These future analyses will use the EDS data, which are not part of Q1.
@@ 7 Conclusions
We have presented the first data release (Q1) of the SPE PF results. We have shown that, if we properly combine the various quality indicators provided by the pipeline, we can reach a success rate above 80% in the redshift range targeted for cosmology ($0.9<z<1.8$). The redshift precision (approximately $10^{-3}$) and accuracy (better than $3\times 10^{-5}$) are also excellent. This is encouraging for the future DR1 and the first cosmological results, but the completeness will still need to be assessed when EDS data will be available and all EWS sources down to $H_{\scriptscriptstyle\rm E}=24$ instead of 22.5 will have been processed.
As would be expected, the results are less good outside of this redshift range, since no strong spectral features are then observable within the NISP red-grism spectral range and these data should be used with caution for non-cosmology science. Future deep-field observations including the blue grism should dramatically improve Euclid’s performance in this regime. In addition, the sample of star candidates selected by SPE PF based on spectroscopy alone is 75% pure. This is very promising for the study of cool stars (Bañados et al., 2025).
Finally, the current results are only a first step and we can expect major improvements in the future. The SIR PF products injected into the SPE PF still contain important artefacts (Euclid Collaboration: Copin et al., 2025), which will be corrected in the next versions of the pipeline. This should dramatically reduce the fraction of spurious redshifts and enable us to use less stringent selection criteria, increasing the size of the high-reliability samples. The SPE PF pipeline itself will also improve. Deep-learning approaches have been developed to estimate the redshift reliability and classify the spectra. We already trained as well as tested such algorithms on simulations and they outperformed the current algorithm. However, we shall need real future EDS data to build a proper training set for the analysis of real EWS data. Finally, the selection of galaxies for cosmological analyses will also benefit from non-spectroscopic information (e.g. morphology or photometry), allowing us to obtain sufficiently pure and complete samples.
@@ Acknowledgements.
MM acknowledges the financial contribution from the grant PRIN-MUR 2022 2022NY2ZRS 001 Optimizing the extraction of cosmological information from Large Scale Structure analysis in view of the next large spectroscopic surveys supported by Next Generation EU and from the grant ASI n. 2024-10-HH.0 Attivit scientifiche per la missione Euclid fase E. The Euclid Consortium acknowledges the European Space Agency and a number of agencies and institutes that have supported the development of Euclid, in particular the Agenzia Spaziale Italiana, the Austrian Forschungsförderungsgesellschaft funded through BMK, the Belgian Science Policy, the Canadian Euclid Consortium, the Deutsches Zentrum für Luft- und Raumfahrt, the DTU Space and the Niels Bohr Institute in Denmark, the French Centre National d’Etudes Spatiales, the Fundação para a Ciência e a Tecnologia, the Hungarian Academy of Sciences, the Ministerio de Ciencia, Innovación y Universidades, the National Aeronautics and Space Administration, the National Astronomical Observatory of Japan, the Netherlandse Onderzoekschool Voor Astronomie, the Norwegian Space Agency, the Research Council of Finland, the Romanian Space Agency, the State Secretariat for Education, Research, and Innovation (SERI) at the Swiss Space Office (SSO), and the United Kingdom Space Agency. A complete and detailed list is available on the Euclid web site (www.euclid-ec.org). This research used data obtained with the Dark Energy Spectroscopic Instrument (DESI). DESI construction and operations is managed by the Lawrence Berkeley National Laboratory. This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of High-Energy Physics, under Contract No. DEAC0205CH11231, and by the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility under the same contract. Additional support for DESI was provided by the U.S. National Science Foundation (NSF), Division of Astronomical Sciences under Contract No. AST-0950945 to the NSFs National Optical-Infrared Astronomy Research Laboratory; the Science and Technology Facilities Council of the United Kingdom; the Gordon and Betty Moore Foundation; the Heising-Simons Foundation; the French Alternative Energies and Atomic Energy Commission (CEA); the National Council of Science and Technology of Mexico (CONACYT); the Ministry of Science and Innovation of Spain (MICINN), and by the DESI Member Institutions: www.desi.lbl.gov/collaborating-institutions. The DESI collaboration is honoured to be permitted to conduct scientific research on Iolkam Duag (Kitt Peak), a mountain with particular significance to the Tohono Oodham Nation. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the U.S. National Science Foundation, the U.S. Department of Energy, or any of the listed funding agencies.
Appendix A Representativeness of the DESI reference sample
In Fig. 9, we present the $H_{\scriptscriptstyle\rm E}$-magnitude distribution of various selections. If we consider the full SPE sample with $H_{\scriptscriptstyle\rm E}$ <<22.5 and $0.9<z_{\rm SPE}<1.6$ (upper left panel), we can observe that number of SPE objects (blue) per magnitude bin increases steeply with increasing magnitude, while the number of objects with both a SPE and a DESI redshift (orange) is significantly flatter than for the full SPE sample above $H_{\scriptscriptstyle\rm E}=20$. Below this magnitude, we can notice a bump peaking around $H_{\scriptscriptstyle\rm E}=18$, as discussed below. Only a small fraction of the $H_{\scriptscriptstyle\rm E}$-faint Euclid galaxies can thus be compared with DESI. This confirms that DESI is not suited to estimate the completeness of the SPE sample. We find a similar behaviour for the high-quality sample described in Sect. 3.3 (top right panel) and the cosmological sample described in Sect. 6.2 (bottom left panel).
We also study the $H_{\scriptscriptstyle\rm E}$-magnitude distribution of objects with a compatible redshift in both surveys ($|z_{\rm SPE}-z_{\rm DESI}|/(1+z_{\rm DESI})<0.003$, red). The comparison of the result with the histogram of all the objects with DESI counterparts provides a direct estimate of the dependence of the fraction of good redshifts with the $H_{\scriptscriptstyle\rm E}$ magnitude. For $H_{\scriptscriptstyle\rm E}$ >>20, the shape of the two histograms is very similar, but with a different normalisation. As expected, the two histograms are very close for the high-purity and cosmological samples (top right and bottom left), since most of the objects in these samples have a correct SPE redshift. This is not the case for the full sample with $H_{\scriptscriptstyle\rm E}$ <<22.5 and $0.9<z_{\rm SPE}<1.6$, for which a large fraction of redshifts are incorrect.
We can clarify the nature of the bump in the $H_{\scriptscriptstyle\rm E}$-magnitude distribution of SPE objects with a DESI redshift below $H_{\scriptscriptstyle\rm E}$ = 20. Most of these sources are low-$z$ DESI sources ($\gtrsim 1/2$ of them are below 0.1 and almost all below 0.9) with a probably incorrect Euclid redshift measured between 0.9 and 1.8. It is thus not surprising that they are bright in the photometry since they are nearby and their low redshift implies that these sources have very few spectral features covered by Euclid spectroscopy and thus unreliable SPE redshifts. In contrast, $H_{\scriptscriptstyle\rm E}<19$ objects actually at $z>0.9$ are rare. Paradoxically, the $H_{\scriptscriptstyle\rm E}$-bright objects measured at $0.9<z_{\rm SPE}<1.6$ are thus less reliable than the fainter ones.
Finally, we computed the fraction of SPE redshifts agreeing with DESI as a function of the $H_{\scriptscriptstyle\rm E}$ magnitude (bottom right panel). Because of the phenomenon discussed above, the fraction is very low up to $H_{\scriptscriptstyle\rm E}$ $\sim$20. Above this limit, the fraction of correct redshifts is mostly stable, except for a slight decrease with increasing magnitude at the faint end. Our estimates of the success rate should thus not be too affected by the fact that only a small fraction of $H_{\scriptscriptstyle\rm E}$-faint objects have a DESI counterpart. We can also note that the cosmological sample has a success rate above 80% in all $H_{\scriptscriptstyle\rm E}$ >>20 bins, which is very encouraging. However, we cannot exclude that the redshift of the sources with a DESI counterpart could be easier to measure, biasing our estimates of the success rate towards optimistic values. For instance, the median $\mathrm{H}\,\alpha$ flux of the cosmological sample is $3.5\times 10^{-16}\,\mathrm{erg\,s}^{-1}\,\mathrm{cm}^{-2}$, while we obtain $4.3\times 10^{-16}\,\mathrm{erg\,s}^{-1}\,\mathrm{cm}^{-2}$ if we restrict our computation to objects with a DESI redshift. This suggests that our DESI reference sample is biased towards brighter objects in $\mathrm{H}\,\alpha$, but also that this bias is rather small.
Figure 9: Figures illustrating the representativeness of DESI as a reference sample. From the top : Number of $0.9<z_{\rm SPE}<1.6$ objects as a function of the $H_{\scriptscriptstyle\rm E}$ magnitude, for the high-quality objects described in Sect. 3.3 and for the cosmological sample described in Sect. 6.2 respectively. The blue, orange, and red histograms are the SPE sample, the sub-sample with a DESI redshift, and the sub-sample with a redshift agreeing with DESI, respectively. Bottom: Fraction of redshifts agreeing with DESI as a function of the $H_{\scriptscriptstyle\rm E}$ magnitude for the $0.9<z_{\rm SPE}<1.6$ SPE sample (black), the sub-sample with a DESI redshift (blue), and the cosmological sample (red).